NvidiaIntely (nvidiaintely):

Which type of radioactive particle(s) has/have been emitted in the following equation?

NvidiaIntely (nvidiaintely):

alpha particle only beta particle only alpha and beta particles alpha and beta particles and gamma radiation

NvidiaIntely (nvidiaintely):

@welshfella @Kevin @brainlesslonelygirl

OpenStudy (kevin):

alpha only

NvidiaIntely (nvidiaintely):

Sure?

OpenStudy (kevin):

222 = 218 + 4 86 = 82 + 2 It should be alpha

NvidiaIntely (nvidiaintely):

Ok